2x+5y=9
4x-7y=-16

Answer by user_dude2008(1862)

2x+5y=9
4x-7y=-16


-4x-10y=-18
4x-7y=-16


-17y=-34
y=2

4x-7y=-16
4x-7(2)=-16
4x-14=-16
4x=-2
x=-1/2

Answers: x=-1/2, y=2